# Assessing Inf luential Observations in Pain Prediction using fMRI Data

### v2

@@ 1 Introduction
To further promote discussion of this topic and gain tangible insights into the pitfalls of ignoring observations exerting disproportionate influence on variable selection, a synthetic numerical example is given in Section 1 of the Supplementary Materials Part I. This illustrates the non-negligible empirical probability of selecting an incorrect submodel based on the full, contaminated dataset. Moreover, this example also explicitly exhibits the adverse impacts of neglecting contaminated observations on the downstream analysis including prediction and model coefficient estimation. Specifically, the mean squared error (MSE) for the predicted responses, the correlation between predicted and true responses, and the MSE for coefficient estimates are all negatively influenced by the presence of influential points. Therefore, given the importance and ubiquity of stochastic variable selection, the possible existence of numerous outliers affecting variable selection especially in fMRI-based pain studies, and the complications caused by ignoring them, it is necessary to identify points influencing the data-dependent choice of a statistical submodel.
@@
This manuscript aims to fill the aforementioned gaps and contribute to influential diagnosis on variable selection in five ways. First, upon characterizing influential points in assessing variable selection, we introduce the notion of sensitivity of model selectors, and draw connections with statistical concepts such as breakdown points and selection consistency. Second, we extend the diagnostic measure of [28] by bridging a missing step in its original theoretical development. Third, a generalized version accommodating universal model selectors is proposed and the corresponding asymptotic distribution determined. The main building blocks in establishing the large sample distribution of our proposed method are the preservation of exchangeability [6] and characterization of a sequence of arbitrarily dependent events by its exchangeable counterpart [13]. Fourth, we propose a novel detection procedure, known as clustering-based MIP (ClusMIP), capable of identifying multiple influential observations. Fifth, we compare our proposal with competing procedures in comprehensive simulation and real data studies. ClusMIP is consistently shown to demonstrate significantly better detection performance than the DF(LASSO). Its application to the scaled LASSO exhibits the highest detection outcome. Consequently, substantial enhancement in prediction and variable selection is observed by removing outliers detected by the ClusMIP. On the other hand, influential points detected by the ClusMIP (applied to MCP) for the fMRI-based pain prediction dataset strengthens the distinction between low and intense pain processing mechanisms. Furthermore, the resulting statistical model is scientifically interpretable and displays the best prediction performance after outlier removal.
@@ 4 Methodology
Let $\text{Y}\in\mathbb{R}$ be the response variable and ${x}_{j}\in\mathbb{R},\,j=1,...,p$, be the predictors. Let $\textbf{Y}=(\text{Y}_{1},...,\text{Y}_{n})^{\top}\in\mathbb{R}^{n}$ be the response vector and $\textbf{X}=(\textbf{X}_{1},...,\textbf{X}_{n})^{\top}\in\mathbb{R}^{n\times p}$, where $\textbf{X}_{i}=(\text{X}_{i1},...,\text{X}_{ip})^{\top}\in\mathbb{R}^{p}$, $i=1,...,n$, be the $n\times p$ design matrix. Then, the $i^{\text{th}}$ observation or point of dataset $\textbf{Z}\coloneqq(\textbf{Y},\textbf{X})$ is $\text{Z}_{i}\coloneqq(\text{Y}_{i},\textbf{X}_{i}^{\top})^{\top}\in\mathbb{R}^{p+1}$, $i=1,..,n$. Moreover, we abbreviate an arbitrary sequence of random variables or real values $a_{1},...,a_{n}$ by $a_{[n]}$.
@@ 4.1.1 High-dimensional Inf luence Measure (HIM)
Based on the linear regression model (1), [37] proposed a non-parametric, model-free and theoretically justified diagnostic measure capable of indirectly assessing the influence of a single observation on variable selection for high-dimensional data. Known as the high-dimensional influence measure (HIM), this metric is formulated on the leave-one-out basis and is designed to quantify the contribution of each individual observation to the marginal correlation between the response variable and all the predictors. An observation with excessively large marginal correlation is deemed influential. To compute the HIM, we first calibrate the influence of the $i^{\text{th}}$ observation on the marginal correlation between the response variable Y and $j^{\text{th}}$ predictor $x_{j}$ by calculating its sample estimate $\widehat{\rho}_{j}^{(i)}=\sum_{k=1,k\neq i}^{n}\big{\{}(\text{X}_{kj}-\widehat{\mu}_{xj}^{(i)})(\text{Y}_{k}-\widehat{\mu}_{y}^{(i)})\big{\}}/\big{\{}(n-1)\widehat{\sigma}_{xj}^{(i)}\widehat{\sigma}_{y}^{(i)}\big{\}}$, $j=1,...,p$ and $i=1,...,n$, where $\widehat{\mu}_{xj}^{(i)}=\sum_{k=1,k\neq i}^{n}\text{X}_{kj}/n$, $\widehat{\mu}_{y}^{(i)}=\sum_{k=1,k\neq i}^{n}\text{Y}_{k}/n$, $\big{(}\widehat{\sigma}_{xj}^{(i)}\big{)}^{2}=\sum_{k=1,k\neq i}^{n}(\text{X}_{kj}-\widehat{\mu}_{xj}^{(i)})^{2}/(n-1)$, and $\big{(}\widehat{\sigma}_{y}^{(i)}\big{)}^{2}=\sum_{k=1,k\neq i}^{n}(\text{Y}_{k}-\widehat{\mu}_{y}^{(i)})^{2}/(n-1)$. Then, upon aggregating and averaging over all $p$ predictors, the HIM for the $i^{\text{th}}$ observation is given by $D_{i}=\sum_{j=1}^{p}\big{(}\widehat{\rho}_{j}-\widehat{\rho}_{j}^{(i)}\big{)}^{2}/p,\ i=1,...,n$, where $\widehat{\rho}_{j}^{(i)}$ is given above and $\widehat{\rho}_{j}$ is its counterpart computed based on the full dataset. Under certain conditions (ibid., Conditions C.1 - C.4, Section 2.3), [37] establish that in the absence of influential points, $n^{2}D_{i}\rightarrow\chi^{2}(1)\ \text{as}\ \text{min}(n,p)\rightarrow\infty,\ i=1,...,n$, where $\chi^{2}(1)$ is the chi-squares distribution with one degree of freedom. Thus, the scaled percentile of $\chi^{2}(1)$ can be used as a threshold to pinpoint influential observations. The authors then demonstrate the potential to extend their framework to GLM, yet remarking that the above asymptotic distributional property may not hold. The effectiveness of this method in assessing influence on variable selection is further validated via simulated studies.
@@ 4.1.2 Difference in Model Selection via the LASSO (DF(LASSO))
In comparison with [37], under (1), [28] formulated the so-called difference in model selection (DF(LASSO)) measure to directly gauge the influence of a single point on the submodel selected by the LASSO. Similar to the HIM, the DF(LASSO) is also based on the leave-one-out scheme and is defined by $\iota_{i}=(\delta_{i}-\mathbb{E}(\delta_{i}))/\sqrt{\mathrm{Var}(\delta_{i})}$, $i=1,...,n$, where
@@ 4.1.3 Generalized Difference in Model Selection Measure (GDF)
To address the first issue regarding the eligibility of estimating $\mathbb{E}(\delta_{i})$ and $\mathrm{Var}(\delta_{i})$ by the corresponding sample mean and variance, we first focus on the finite-sample property of $\delta_{[n]}$ and bridge the gap between this sequence and the characteristic of the original dataset. With Theorem 1 reported below, we show that $\delta_{[n]}$ is exchangeable and thus identically distributed. The proof of Theorem 1 is given in Section 2 of the Supplementary Materials Part I.
@@
From Theorem 1, we next provide two relevant results of an exchangeable sequence of random variables in Corollary 1, which are the consistency of the empirical cumulative distribution function (ECDF) and strong law of large numbers (SLLN). The combined statements further strengthen the theoretical basis for estimating $\mathbb{E}(\delta_{i})$ and $\mathrm{Var}(\delta_{i})$ by the sample mean and variance. Its proof is given in Section 3 of the Supplementary Materials Part I.
@@
[…]
@@
[…]
@@
To verify the CLT for $\tau_{i}$, we bridge the theoretical hiatus in [28] for generic model selectors through the work of [13], which bypasses the challenge in handling the intractable dependence among its summands $\xi_{i,j}$. Specifically, it is shown [ibid., Theorem 1] that for an arbitrary sequence of events, there always exists a sequence of exchangeable events such that the respective sums are equally distributed in finite sample. This implies that $\tau_{i}$ in (3), the sum of dependent indicators $\xi_{i,j},j=1,...,p$, is identically distributed as the sum of a sequence of exchangeable Bernoulli random variables. Toward that end, the regular [19] and empirical [15] CLT for an exchangeable sequence of random variables collectively facilitate the derivation of the following assertion, where its proof is available in Section 4 of the Supplementary Materials Part I:
@@
Let $\tau_{i}$ be the GDF measure defined as in (3). Suppose that the conditions (A.1) and (A.2) (see Section 4 of the Supplementary Materials Part I) hold. Then,
@@
While we essentially have summation of Bernoulli variables, we have presented the conditions in rather general form to extend the applicability of the result. Here, condition (A.1) controls the tail behavior of the exchangeable summand of $\tau_{i}$ by requiring that it is not heavy tailed, which is a weak version of the existence of the first moment. Condition (A.2) is parallel to the classical Lindeberg’s condition in establishing the CLT, which requires that the contribution of each individual summand to the variance of $\tau_{i}$ is sufficiently small. In fact, the existence of a uniform bound on the second moment of the summand serves as a sufficient (not necessary) condition for (A.1) and (A.2). In our case, each individual summand is Bernoulli so that both conditions are fulfilled. From (4), at a given nominal level $\alpha\in(0,1)$, the $100\cdot(1-\alpha/2)^{\text{th}}$ percentile of $N(0,1)$ is a valid threshold such that a point with an absolute scale of the standardized $\tau_{i}$ exceeding that cut-off is deemed influential. Moreover, to test the two-sided hypothesis that the $i^{\text{th}}$ point is clean, the $p$-value is $2\cdot\mathbb{P}(N(0,1)>(\tau_{i}-\bar{\tau})/\sigma_{\tau})$. A framework for detecting a single influential point via the GDF metric is given in Algorithm GDF in Section 5 of the Supplementary Materials Part I.
@@ 4.2.1 Multiple Inf luential Point (MIP) Detection Procedure
To have a proper understanding of this procedure, we first explain the philosophy underlying the RGD algorithm, which is provided in Section 6 of the Supplementary Materials Part I. The RDG scheme centers on seeking and subsequently attaching a clean subset of data to every single point, of which refinement is conducted based on the combined dataset. To be more precise, for each observation $i$, $m$ samples indexed by $A_{1},...,A_{m}$ are drawn uniformly and randomly from the remaining portion of the data. When $m$ is sufficiently large, there is a high probability that the subset $\{\text{Z}_{r}:r\in A_{j}\}$ is clean for some $j\in\{1,...,m\}$. Thus, if $i^{\text{th}}$ observation is contaminated, it would be the sole influential point in the merged subset $\{\text{Z}_{r}:r\in A_{j}\cup\{i\}\}$. Thus, further assessment can be carried out on this combined dataset.
Toward that end, for each $i$, two HIM-based statistics $\text{T}_{\text{min},i,m}$ and $\text{T}_{\text{max},i,m}$ are then constructed, which are respectively defined by $\text{T}_{\text{min},i,m}=\arg\!\min_{1\leq j\leq m}n_{\text{sub}}^{2}D_{j,i}$ and $\text{T}_{\text{max},i,m}=\mathop{\mathrm{argmax}}_{1\leq j\leq m}n_{\text{sub}}^{2}D_{j,i}$, where $D_{j,i}=\sum_{k=1}^{p}\big{(}\widehat{\rho}_{k}^{(A_{j}\cup\{i\})}-\widehat{\rho}_{k}^{(A_{j}}\big{)}^{2}/p$. Here, $\widehat{\rho}_{k}^{(A_{j}\cup\{i\})}$ and $\widehat{\rho}_{k}^{(A_{j})}$ are the sample marginal correlations between the response Y and $k^{\text{th}}$ predictor $x_{k}$ based on $\{\text{Z}_{r}:r\in\{A_{j}\cup\{i\}\}\}$ and $\{\text{Z}_{r}:r\in A_{j}\}$. The authors then establish that they are both asymptotically $\chi^{2}(1)$ and indeed complement each other: $\text{T}_{\text{min},i,m}$ tends to select conspicuously influential points. Thus, it is effective in overcoming the swamping effect. However, it may be conservative so that true influential points may be neglected. In contrast, $\text{T}_{\text{max},i,m}$ is powerful to counter the masking effect. Yet, it may be inordinately aggressive so that clean ones may be falsely chosen. Therefore, combining the strengths of both $\text{T}_{\text{min},i,m}$ and $\text{T}_{\text{max},i,m}$ leads to the MIP detection procedure provided in Section 7 of the Supplementary Materials Part I. In practice, the MIP algorithm is implemented in the $\tt{R}$ package $\tt{MIP}$.

### v3

@@ 1 Introduction
To further promote discussion of this topic and gain tangible insights into the pitfalls of ignoring observations exerting disproportionate influence on variable selection, a synthetic numerical example is given in Section 1 of the Supplementary Materials Part I (Zhang, Asgharian and Lindquist (2026)). This illustrates the non-negligible empirical probability of selecting an incorrect submodel based on the full, contaminated dataset. Moreover, this example also explicitly exhibits the adverse impacts of neglecting contaminated observations on the downstream analysis including prediction and model coefficient estimation. Specifically, the mean squared error (MSE) for the predicted responses, the correlation between predicted and true responses, and the MSE for coefficient estimates are all negatively influenced by the presence of influential points. Therefore, given the importance and ubiquity of stochastic variable selection, the possible existence of numerous outliers affecting variable selection especially in fMRI-based pain studies, and the complications caused by ignoring them, it is necessary to identify points influencing the data-dependent choice of a statistical submodel.
@@
This manuscript aims to fill the aforementioned gaps and contribute to influential diagnosis on variable selection in five ways. First, upon characterizing influential points in assessing variable selection, we introduce the notion of sensitivity of model selectors, and draw connections with statistical concepts such as breakdown points and selection consistency. Second, we extend the diagnostic measure of [28] by bridging a missing step in its original theoretical development. Third, a generalized version accommodating universal model selectors is proposed and the corresponding asymptotic distribution determined. The main building blocks in establishing the large sample distribution of our proposed method are the preservation of exchangeability [6] and characterization of a sequence of arbitrarily dependent events by its exchangeable counterpart [13]. Fourth, we propose a novel detection procedure, known as clustering-based MIP (ClusMIP), capable of identifying multiple influential observations. Fifth, we compare our proposal with competing procedures in comprehensive simulation and real data studies. ClusMIP is consistently shown to demonstrate significantly better detection performance than DF(LASSO). Its application to the scaled LASSO exhibits the highest detection outcome (Panels (A) of Figures 1 and 2). This implies that the (scaled) LASSO is the most sensitive model selector to influential points with regards to variable selection, and therefore serves as an appropriate candidate to use together with the ClusMIP procedure for outlier identification purposes. As such, our simulations offers insights for practitioners into the sensitivity of different model selectors to outliers on variable selection performance under contamination. In addition, substantial enhancement in prediction and variable selection is observed by removing outliers detected by the ClusMIP. On the other hand, influential points detected by the ClusMIP (applied to MCP) for the fMRI-based pain prediction dataset strengthens the distinction between low and intense pain processing mechanisms. Furthermore, the resulting statistical model is scientifically interpretable and displays the best prediction performance after outlier removal.
@@ 4 Methodology
Let $\text{Y}\in\mathbb{R}$ be the response variable and ${x}_{j}\in\mathbb{R},\,j=1,...,p$, be the predictors. Let $\textbf{Y}=(\text{Y}_{1},...,\text{Y}_{n})^{\top}\in\mathbb{R}^{n}$ be the response vector and $\textbf{X}=(\textbf{X}_{1},...,\textbf{X}_{n})^{\top}\in\mathbb{R}^{n\times p}$, where $\textbf{X}_{i}=(\text{X}_{i1},...,\text{X}_{ip})^{\top}\in\mathbb{R}^{p}$, i=1,,ni=1,...,n, be the $n\times p$ design matrix. Then, the $i^{\text{th}}$ observation or point of dataset $\textbf{Z}\coloneqq(\textbf{Y},\textbf{X})$ is $\text{Z}_{i}\coloneqq(\text{Y}_{i},\textbf{X}_{i}^{\top})^{\top}\in\mathbb{R}^{p+1}$, $i=1,..,n$. Moreover, we abbreviate an arbitrary sequence of random variables or real values $a_{1},...,a_{n}$ by $a_{[n]}$.
@@ 4.1.1 High-dimensional Influence Measure (HIM)
Based on the linear regression model (1), [37] proposed a non-parametric, model-free and theoretically justified diagnostic measure capable of indirectly assessing the influence of a single observation on variable selection for high-dimensional data. Known as the high-dimensional influence measure (HIM), this metric is formulated on the leave-one-out basis and is designed to quantify the contribution of each individual observation to the marginal correlation between the response variable and all the predictors. An observation with excessively large marginal correlation is deemed influential. To compute the HIM, we first calibrate the influence of the $i^{\text{th}}$ observation on the marginal correlation between the response variable Y and $j^{\text{th}}$ predictor $x_{j}$ by calculating its sample estimate $\widehat{\rho}_{j}^{(i)}=\sum_{k=1,k\neq i}^{n}\big\{(\text{X}_{kj}-\widehat{\mu}_{xj}^{(i)})(\text{Y}_{k}-\widehat{\mu}_{y}^{(i)})\big\}/\big\{(n-1)\widehat{\sigma}_{xj}^{(i)}\widehat{\sigma}_{y}^{(i)}\big\}$, j=1,,pj=1,...,p and i=1,,ni=1,...,n, where $\widehat{\mu}_{xj}^{(i)}=\sum_{k=1,k\neq i}^{n}\text{X}_{kj}/n$, $\widehat{\mu}_{y}^{(i)}=\sum_{k=1,k\neq i}^{n}\text{Y}_{k}/n$, $\big(\widehat{\sigma}_{xj}^{(i)}\big)^{2}=\sum_{k=1,k\neq i}^{n}(\text{X}_{kj}-\widehat{\mu}_{xj}^{(i)})^{2}/(n-1)$, and $\big(\widehat{\sigma}_{y}^{(i)}\big)^{2}=\sum_{k=1,k\neq i}^{n}(\text{Y}_{k}-\widehat{\mu}_{y}^{(i)})^{2}/(n-1)$. Then, upon aggregating and averaging over all $p$ predictors, the HIM for the $i^{\text{th}}$ observation is given by $D_{i}=\sum_{j=1}^{p}\big(\widehat{\rho}_{j}-\widehat{\rho}_{j}^{(i)}\big)^{2}/p,\ i=1,...,n$, where $\widehat{\rho}_{j}^{(i)}$ is given above and $\widehat{\rho}_{j}$ is its counterpart computed based on the full dataset. Under certain conditions (ibid., Conditions C.1 - C.4, Section 2.3), [37] establish that in the absence of influential points, $n^{2}D_{i}\rightarrow\chi^{2}(1)\ \text{as}\ \text{min}(n,p)\rightarrow\infty,\ i=1,...,n$, where $\chi^{2}(1)$ is the chi-squares distribution with one degree of freedom. Thus, the scaled percentile of $\chi^{2}(1)$ can be used as a threshold to pinpoint influential observations. The authors then demonstrate the potential to extend their framework to GLM, yet remarking that the above asymptotic distributional property may not hold. The effectiveness of this method in assessing influence on variable selection is further validated via simulated studies.
@@ 4.1.2 Difference in Model Selection via the LASSO (DF(LASSO))
In comparison with [37], under (1), [28] formulated the so-called difference in model selection (DF(LASSO)) measure to directly gauge the influence of a single point on the submodel selected by the LASSO. Similar to the HIM, the DF(LASSO) is also based on the leave-one-out scheme and is defined by $\iota_{i}=(\delta_{i}-\mathbb{E}(\delta_{i}))/\sqrt{\mathrm{Var}(\delta_{i})}$, i=1,,ni=1,...,n, where
@@ 4.1.3 Generalized Difference in Model Selection Measure (GDF)
To address the first issue regarding the eligibility of estimating $\mathbb{E}(\delta_{i})$ and $\mathrm{Var}(\delta_{i})$ by the corresponding sample mean and variance, we first focus on the finite-sample property of $\delta_{[n]}$ and bridge the gap between this sequence and the characteristic of the original dataset. With Theorem 1 reported below, we show that $\delta_{[n]}$ is exchangeable and thus identically distributed. The proof of Theorem 1 is given in Section 2 of the Supplementary Materials Part I (Zhang, Asgharian and Lindquist (2026)).
@@
From Theorem 1, we next provide two relevant results of an exchangeable sequence of random variables in Corollary 1, which are the consistency of the empirical cumulative distribution function (ECDF) and strong law of large numbers (SLLN). The combined statements further strengthen the theoretical basis for estimating $\mathbb{E}(\delta_{i})$ and $\mathrm{Var}(\delta_{i})$ by the sample mean and variance. Its proof is given in Section 3 of the Supplementary Materials Part I (Zhang, Asgharian and Lindquist (2026)).
@@
    Uniform Convergence of ECDF: $\sup_{x\in\mathbb{R}}|\widehat{F}_{n}-F|\xrightarrow{\text{a.s.}}0$, where $\widehat{F}_{n}(x)=\sum_{i=1}^{n}\mathbbm{1}(\delta_{i}\leq x)$, $x\in\mathbb{R}$ and $F$ is the true cumulative distribution function (CDF) of $\delta_{[n]}$. In particular, the discontinuity of the limiting CDF $F$ is permitted.
@@
    SLLN: $\frac{1}{n}\sum_{i=1}^{n}\delta_{i}\xrightarrow{\text{a.s.}}\mathbb{E}(\delta_{i}|\mathcal{F}_{\infty})$ and $\frac{1}{n}\sum_{i=1}^{n}(\delta_{i}-\bar{\delta})^{2}\xrightarrow{\text{a.s.}}\mathrm{Var}(\delta_{i}|\mathcal{F}_{\infty})$.
@@
To verify the CLT for $\tau_{i}$, we bridge the theoretical hiatus in [28] for generic model selectors through the work of [13], which bypasses the challenge in handling the intractable dependence among its summands $\xi_{i,j}$. Specifically, it is shown [ibid., Theorem 1] that for an arbitrary sequence of events, there always exists a sequence of exchangeable events such that the respective sums are equally distributed in finite sample. This implies that $\tau_{i}$ in (3), the sum of dependent indicators $\xi_{i,j},j=1,...,p$, is identically distributed as the sum of a sequence of exchangeable Bernoulli random variables. Toward that end, the regular [19] and empirical [15] CLT for an exchangeable sequence of random variables collectively facilitate the derivation of the following assertion, where its proof is available in Section 4 of the Supplementary Materials Part I (Zhang, Asgharian and Lindquist (2026)):
@@
Let $\tau_{i}$ be the GDF measure defined as in (3). Suppose that the conditions (A.1) and (A.2) (see Section 4 of the Supplementary Materials Part I (Zhang, Asgharian and Lindquist (2026))) hold. Then,
@@
While we essentially have summation of Bernoulli variables, we have presented the conditions in rather general form to extend the applicability of the result. Here, condition (A.1) controls the tail behavior of the exchangeable summand of $\tau_{i}$ by requiring that it is not heavy tailed, which is a weak version of the existence of the first moment. Condition (A.2) is parallel to the classical Lindeberg’s condition in establishing the CLT, which requires that the contribution of each individual summand to the variance of $\tau_{i}$ is sufficiently small. In fact, the existence of a uniform bound on the second moment of the summand serves as a sufficient (not necessary) condition for (A.1) and (A.2). In our case, each individual summand is Bernoulli so that both conditions are fulfilled. From (4), at a given nominal level $\alpha\in(0,1)$, the $100\cdot(1-\alpha/2)^{\text{th}}$ percentile of $N(0,1)$ is a valid threshold such that a point with an absolute scale of the standardized $\tau_{i}$ exceeding that cut-off is deemed influential. Moreover, to test the two-sided hypothesis that the $i^{\text{th}}$ point is clean, the $p$-value is $2\cdot\mathbb{P}(N(0,1)>(\tau_{i}-\bar{\tau})/\sigma_{\tau})$. A framework for detecting a single influential point via the GDF metric is given in Algorithm GDF in Section 5 of the Supplementary Materials Part I (Zhang, Asgharian and Lindquist (2026)).
@@ 4.2.1 Multiple Influential Point (MIP) Detection Procedure
To have a proper understanding of this procedure, we first explain the philosophy underlying the RGD algorithm, which is provided in Section 6 of the Supplementary Materials Part I (Zhang, Asgharian and Lindquist (2026)). The RDG scheme centers on seeking and subsequently attaching a clean subset of data to every single point, of which refinement is conducted based on the combined dataset. To be more precise, for each observation $i$, $m$ samples indexed by $A_{1},...,A_{m}$ are drawn uniformly and randomly from the remaining portion of the data. When $m$ is sufficiently large, there is a high probability that the subset $\{\text{Z}_{r}:r\in A_{j}\}$ is clean for some $j\in\{1,...,m\}$. Thus, if $i^{\text{th}}$ observation is contaminated, it would be the sole influential point in the merged subset $\{\text{Z}_{r}:r\in A_{j}\cup\{i\}\}$. Thus, further assessment can be carried out on this combined dataset.
Toward that end, for each $i$, two HIM-based statistics $\text{T}_{\text{min},i,m}$ and $\text{T}_{\text{max},i,m}$ are then constructed, which are respectively defined by $\text{T}_{\text{min},i,m}=\arg\!\min_{1\leq j\leq m}n_{\text{sub}}^{2}D_{j,i}$ and $\text{T}_{\text{max},i,m}=\mathop{\mathrm{argmax}}_{1\leq j\leq m}n_{\text{sub}}^{2}D_{j,i}$, where $D_{j,i}=\sum_{k=1}^{p}\big(\widehat{\rho}_{k}^{(A_{j}\cup\{i\})}-\widehat{\rho}_{k}^{(A_{j}}\big)^{2}/p$. Here, $\widehat{\rho}_{k}^{(A_{j}\cup\{i\})}$ and $\widehat{\rho}_{k}^{(A_{j})}$ are the sample marginal correlations between the response Y and $k^{\text{th}}$ predictor $x_{k}$ based on $\{\text{Z}_{r}:r\in\{A_{j}\cup\{i\}\}\}$ and $\{\text{Z}_{r}:r\in A_{j}\}$. The authors then establish that they are both asymptotically $\chi^{2}(1)$ and indeed complement each other: $\text{T}_{\text{min},i,m}$ tends to select conspicuously influential points. Thus, it is effective in overcoming the swamping effect. However, it may be conservative so that true influential points may be neglected. In contrast, $\text{T}_{\text{max},i,m}$ is powerful to counter the masking effect. Yet, it may be inordinately aggressive so that clean ones may be falsely chosen. Therefore, combining the strengths of both $\text{T}_{\text{min},i,m}$ and $\text{T}_{\text{max},i,m}$ leads to the MIP detection procedure provided in Section 7 of the Supplementary Materials Part I (Zhang, Asgharian and Lindquist (2026)). In practice, the MIP algorithm is implemented in the $\tt{R}$ package $\tt{MIP}$.